OpenStudy (kobeni-chan):

Will Medal and Fan, please help? Trig Question: Need to find two vertical asymptotes for f(x) = - tan 4x? I heard that you need to find values when cos(x)=0, but I don't know how to find x...

OpenStudy (kobeni-chan):

@sweetburger can you help please?

OpenStudy (crabbyoldgamer):

Do you know the unit circle?

OpenStudy (michele_laino):

hint: the period of the function \(x \rightarrow -\tan x\) is \(T_0=\pi\), and its vertical asymptotes are located at \(x= \pm \pi/2, \pm 3\pi/2,...\) Now the period of the function \(x \rightarrow -\tan (4x)\) is \(T_1=T_0/4\), so every value of \(x\) at which there is an asymptote for the function \(x \rightarrow -\tan x\) is reduced by a factor \(4\)

OpenStudy (anonymous):

Refer to the attached plot.